# Large language model psychometrics: A systematic review of evaluation, validation, and enhancement

### v2

@@ 4.2.1 Heuristics and biases
Heuristics and biases are mental shortcuts that simplify decision-making but can introduce systematic errors [Tversky and Kahneman, 1974]. Recent research employs psychometric tools to evaluate rationality and biases in LLM outputs and offer theoretical explanations for these biases.
Early studies, such as Binz and Schulz [2023], found that GPT-3 performs comparably to humans on cognitive ability tests, exhibiting biases like framing and anchoring. More advanced models, such as GPT-4, show fewer System 1 (intuitive) errors and more System 2 (deliberative) reasoning, though biases persist [Hagendorff et al., 2023]. Large-scale evaluations now cover a wide range of biases—including conjunction fallacy, unwarranted beliefs, and loss aversion—using synthetic scenarios and multi-agent dialogue [Malberg et al., 2024, Xie et al., 2024b]. Multimodal models are also being assessed for similar patterns [Schulze Buschoff et al., 2025].
To explain these findings, researchers draw on dual-process theory, linking LLM reasoning to intuitive and deliberative modes [Hagendorff et al., 2023, Yax et al., 2024]. Other frameworks highlight LLMs’ lack of metacognitive abilities, leading to “metacognitive myopia”—the inability to monitor or control reasoning processes, which perpetuates systematic biases [Scholten et al., 2024]. Additional perspectives use cognitive dissonance and elaboration likelihood theories to interpret inconsistencies [Sundaram and Alwar, 2024], while mechanistic approaches leverage influence graphs and Shapley values to trace the origins of biases in model training [Shaikh et al., 2024].
@@
By administering psychometrics to LLMs, studies consistently find that LLMs exhibit cognitive biases superficially similar to humans, such as anchoring, framing, and the conjunction fallacy. Large-scale tests, such as [Echterhoff et al., 2024, Malberg et al., 2024, Xie et al., 2024b], systematically categorize these biases and enable scalable evaluations. Some note that newer, larger, chain-of-thought-enabled models exhibit improved reasoning and bias mitigation [Tang and Kejriwal, 2024, Hagendorff et al., 2023], while others argue that increasing model complexity without deliberate bias mitigation strategies can amplify existing biases [Kumar et al., 2024]. In addition, researchers explore mechanistic interpretability [Shaikh et al., 2024] and present theoretical explanations from cognitive psychology and reasoning theories [Hagendorff et al., 2023, Yax et al., 2024, Scholten et al., 2024]. While LLMs exhibit dual-process reasoning dynamics reminiscent of human cognition [Hagendorff et al., 2023], detailed analysis reveals differences between LLM and human reasoning [Yax et al., 2024].
@@ 4.2.2 Social interactions
ToM is the ability to attribute mental states such as beliefs, intentions, and knowledge to others [Premack and Woodruff, 1978]. Advanced LLMs can simulate ToM-like reasoning under certain conditions, which prompts questions about how these behaviors arise and how robustly they generalize. Foundational studies apply classic psychometrics—false belief tasks—to evaluate ToM in LLMs. [Kosinski, 2023a, b] report that GPT-3.5 and GPT-4 perform at levels matching or exceeding those of children in structured ToM tasks, suggesting the spontaneous emergence of ToM-like behavior as models scale in size. Through mechanistic interpretation, Jamali et al. [2023] offer evidence of parallels between the LLM embeddings and neurons in the human brain.
However, these claims have been challenged. Ullman [2023], Shapira et al. [2023a], Holterman and Deemter [2023] show that minor task changes often cause LLMs’ performance to drop, suggesting reliance on brittle heuristics rather than true semantic understanding. To address this, researchers have developed more robust benchmarks and evaluation protocols featuring procedurally generated, higher-order, or broader ToM tasks [Gandhi et al., 2023, Chen et al., 2024d, He et al., 2023, Xu et al., 2024a, Shinoda et al., 2025]. These benchmarks frequently reveal significant declines in performance on complex tasks, such as 6th-order belief attribution [Street et al., 2024] or faux pas detection [Strachan et al., 2024a].
Some argue these failures stem from a lack of general commonsense reasoning rather than an inability to represent mental states [Pi et al., 2024]. Riemer et al. [2025] further suggest distinguishing between literal and functional ToM, noting that current benchmarks inadequately assess functional ToM. While newer models continue to improve on structured ToM tasks [Strachan et al., 2024a, Street et al., 2024], they remain inconsistent in open-ended, adversarial, or pragmatic reasoning scenarios [Sclar et al., 2024, Yu et al., 2025, Nickel et al., 2024]. Comparisons with human baselines show that LLMs can approximate ToM behavior in narrow contexts but still fall short of general human-like social cognition [Strachan et al., 2024a, van Duijn et al., 2023, Jones et al., 2024]. Recent work has also extended ToM evaluation to multimodal settings [Chen et al., 2024c, Strachan et al., 2024b, Jin et al., 2024a] and multi-agent interactions [Li et al., 2023b].
@@
LLMs exhibit psychometrically measurable ToM-like reasoning, especially when appropriately prompted or structured, but current evidence suggests these capabilities depend on surface-level linguistic cues and lack robustness. The proficiency of LLMs in ToM remains a contentious issue. Interested readers are referred to recent comprehensive reviews on ToM in LLMs [Sarıtaş et al., 2025, Ma et al., 2023, Mao et al., 2024b].
@@
EI is the subset of social intelligence that involves the ability to monitor one’s own and others’ feelings and emotions, to discriminate among them and to use this information to guide one’s thinking and actions [Salovey and Mayer, 1990].
Recent work introduces benchmarks to assess EI in LLMs using structured, theory-driven tasks. Tools like EmoBench [Sabour et al., 2024] and EQ-Bench [Paech, 2023] operationalize constructs such as emotional understanding and regulation, often referencing frameworks like Mayer-Salovey-Caruso Emotional Intelligence Test. Other studies adapt established psychometric instruments, including Situational Evaluation of Complex Emotional Understanding (SECEU) [Wang et al., 2023b], the Levels of Emotional Awareness Scale (LEAS) [Elyoseph et al., 2023], the Toronto Alexithymia Scale (TAS-20) and Empathy Quotient (EQ-60) [Patel and Fan, 2023], among others [Huang et al., 2024b]. Advanced models like GPT-4 often match or surpass human baselines in emotional awareness and understanding [Wang et al., 2023b, Elyoseph et al., 2023, Patel and Fan, 2023], though they lack deep reflexive analysis of emotional experience and motivation [Vzorinab et al., 2024]. Recent benchmarks also extend EI evaluation to multimodal settings [Hu et al., 2025].
@@
Advanced LLMs generally perform on par with or better than humans in tasks of EI [Elyoseph et al., 2023, Patel and Fan, 2023], despite evident limitations in several areas, such as displaying artificial or mechanical patterns when expressing empathy [Lee et al., 2024f], the deep reflexive analysis of emotional experiences [Vzorinab et al., 2024], and the misalignment with human emotional behaviors [Huang et al., 2023b].
@@
SI is the ability to understand and manage people [Thorndike and Stein, 1937]. SI in LLMs determines how well these models interpret and respond to social situations. Psychometric tools applied to LLMs yield mixed results. Mittelstädt et al. [2024] use Situational Judgment Tests (SJTs) and find some LLMs outperform humans in expert-rated social appropriateness. However, LLMs still struggle with implicit social understanding: Shapira et al. [2023b] show LLMs perform poorly on faux pas tests, and Xu et al. [2024b] find superficial friendliness often leads to errors in the Situational Evaluation of SI (SESI). The AgentSense benchmark further highlights LLMs’ limitations in complex social interactions, especially regarding high-level needs and private information [Mou et al., 2024].
Recent work expands to multi-agent environments. SOTOPIA [Zhou et al., 2024b] and SOTOPIA-$\pi$ [Wang et al., 2024d] simulate cooperative and competitive social contexts, while the STSS benchmark [Wang et al., 2024a] evaluates SI through task-oriented simulations focused on outcomes and goals. The CogMir framework [Liu et al., 2024c] finds LLMs and humans are similarly consistent in irrational and prosocial decisions under uncertainty. DeSIQ [Guo et al., 2023a] extends SI evaluation to multimodal settings. Critically, Kovač et al. [2024] argue current benchmarks lack developmental grounding and propose the SocialAI school as a more comprehensive framework for studying LLM-based SI.
@@
LLMs show measurable competence in rule-based, socially appropriate behavior, particularly in structured environments. They perform well on: 1) following predefined social norms [Mittelstädt et al., 2024]; 2) completing interactional goals in multi-agent settings [Zhou et al., 2024b, Wang et al., 2024d]; 3) replicating human-like prosocial decisions [Liu et al., 2024c]; and 4) being persuasive, even better than human persuaders [Schoenegger et al., 2025]. Weaker performance is noted in tasks where superficial friendliness causes errors [Xu et al., 2024b], understanding high-level growth needs is essential [Mou et al., 2024], and LLMs must implicitly describe social situations [Shapira et al., 2023b].
@@ 4.2.3 Psychology of language
Psycholinguistics, a subfield of psychology, explores how humans comprehend, generate, and acquire language [Carroll, 1986]. Insights from this discipline help evaluate how LLMs process language and mirror human linguistic features.
@@
LLM language comprehension is evaluated across multiple linguistic levels—sound, word, syntax, meaning, and discourse. Duan et al. [2024a] introduce a benchmark of 10 psycholinguistic tests to assess these comprehensively. Many studies focus on specific aspects, such as syntactic and semantic processing [Wang et al., 2024b, Arehalli et al., 2022, Wolfman et al., 2024, Wilcox et al., 2021]. Surprisal, a measure of word predictability, is widely used to model processing difficulty in syntactic ambiguities and hierarchical structures [Wilcox et al., 2021, Arehalli et al., 2022, Li et al., 2024a]. Results are mixed: for example, GPT-2-XL shows human-like competence in sound-gender association and implicit causality, but not in sound-shape association [Duan et al., 2024b]. Other evaluation tasks include grammaticality judgment [Qiu et al., 2024, Ide et al., 2024], pragmatic inference [Bojić et al., 2023, Hu et al., 2023, Ruis et al., 2023], argument role processing [Lee et al., 2024b], and discourse comprehension [Duan et al., 2024a]. Advanced models like GPT-4 often match or surpass humans in pragmatic reasoning [Bojić et al., 2023] and grammaticality judgment [Dentella et al., 2024], though performance varies with prompt format [Hu and Levy, 2023] and implicature understanding remains limited [Ruis et al., 2023].
@@
A rich line of research evaluates the creativity of LLM language generation [Tang and Kejriwal, 2024, Stevenson et al., 2022, Hubert et al., 2023, Chakrabarty et al., 2024, Orwig et al., 2024, Lee and Chung, 2024, Boussioux et al., 2024, Bellemare-Pepin et al., 2024], using tests like Guilford’s Alternative Uses Test (AUT) [Guilford et al., 1978] and the Torrance Test of Creative Thinking (TTCT) [Torrance, 1966]. Early models such as GPT-3 lack originality and novelty, while advanced LLMs like GPT-4 surpass the human average in creativity. Notably, Tang and Kejriwal [2024] find LLMs underperform in divergent creativity (e.g., novel uses for familiar objects) but can rival humans in open-ended creative writing. Cai et al. [2024] show that ChatGPT produces human-like responses in most of 12 psycholinguistic tests, but AI-generated analogies often lack human-like linguistic properties [Seals and Shalin, 2023]. LLM-generated stories tend to be positive and lack suspense or narrative diversity compared to human writing [Tian et al., 2024]. Linguistic profiling has also been used to characterize the task-specific language abilities of LLMs [Miaschi et al., 2024].
@@
Work on developmental plausibility examines whether LLMs replicate stages of language learning analogous to those in children [Steuer et al., 2023, Shah et al., 2024]. It is shown that regardless of model size, the developmental trajectories of pretrained language models consistently exhibit a window of maximal alignment with human cognitive development [Shah et al., 2024]. Another study by Frank [2023b] draws inspiration from human language development to explain the data inefficiency of LLMs; the relative efficiency of human language acquisition is possibly due to pre-existing conceptual knowledge, multimodal grounding, and the interactive, social nature of their input.
@@
Early foundations that evaluate language models based on BERT and LSTM link computational linguistics and psycholinguistic mechanisms [Ettinger, 2020, Futrell et al., 2019, Arehalli and Linzen, 2020]; these models generally fall short in diverse psycholinguistic tasks. Recent studies extend evaluations to LLMs and broader tasks [Duan et al., 2024a], where advanced LLMs are shown to surpass humans on tasks such as pragmatic reasoning [Bojić et al., 2023] and creative writing [Tang and Kejriwal, 2024]. Some deficiencies still exist, such as limited implicature understanding [Ruis et al., 2023], prompt-sensitive linguistic competence [Hu and Levy, 2023], and lack of human-like psycholinguistic properties [Seals and Shalin, 2023, Tian et al., 2024]. In addition, mixed results are found in the alignment between LLMs and human linguistic cognition [Duan et al., 2024b, Wolfman et al., 2024] and language acquisition [Shah et al., 2024, Frank, 2023b].

### v3

@@ 4.2.1 Heuristics and biases
Heuristics and biases are mental shortcuts that simplify decision-making but can introduce systematic errors [tversky1974judgment]. Recent research employs psychometric tools to evaluate rationality and biases in LLM outputs and offer theoretical explanations for these biases.
Early studies, such as binz2023using, found that GPT-3 performs comparably to humans on cognitive ability tests, exhibiting biases like framing and anchoring. More advanced models, such as GPT-4, show fewer System 1 (intuitive) errors and more System 2 (deliberative) reasoning, though biases persist [hagendorff2023reasoning_bias]. Large-scale evaluations now cover a wide range of biases (e.g., conjunction fallacy, unwarranted beliefs, and loss aversion) using synthetic scenarios and multi-agent dialogue [malberg2024comprehensive_cognitive_biases, xie2024mindscope]. Multimodal models are also being assessed for similar patterns [schulze2025visual].
To explain these findings, researchers draw on dual-process theory, linking LLM reasoning to intuitive and deliberative modes [hagendorff2023reasoning_bias, yax2024studying]. Other frameworks highlight LLMs’ lack of metacognitive abilities, i.e., the inability to monitor or control reasoning processes. Such inability perpetuates systematic biases in LLMs [scholten2024metacognitive]. Additional perspectives use cognitive dissonance and elaboration likelihood theories to interpret inconsistencies [sundaram2024can], while mechanistic approaches leverage influence graphs and Shapley values to trace the origins of biases in model training [shaikh2024cbeval].
@@
By administering psychometrics to LLMs, studies consistently find that LLMs exhibit cognitive biases superficially similar to humans, such as anchoring, framing, and the conjunction fallacy. Large-scale tests, such as [echterhoff2024cognitive, malberg2024comprehensive_cognitive_biases, xie2024mindscope], systematically categorize these biases and enable scalable evaluations. Some note that newer, larger, chain-of-thought-enabled models exhibit improved reasoning and bias mitigation [tang2024humanlike, hagendorff2023reasoning_bias], while others argue that increasing model complexity without deliberate bias mitigation strategies can amplify existing biases [kumar2024investigating]. In addition, researchers explore mechanistic interpretability [shaikh2024cbeval] and present theoretical explanations from cognitive psychology and reasoning theories [hagendorff2023reasoning_bias, yax2024studying, scholten2024metacognitive]. While LLMs exhibit dual-process reasoning dynamics reminiscent of human cognition [hagendorff2023reasoning_bias], detailed analysis reveals differences between LLM and human reasoning [yax2024studying].
@@ 4.2.2 Social interactions
ToM is the ability to attribute mental states such as beliefs, intentions, and knowledge to others [premack1978tom]. Advanced LLMs can simulate ToM-like reasoning under certain conditions, which prompts questions about how these behaviors arise and how robustly they generalize. Foundational studies apply classic psychometrics (false belief tasks) to evaluate ToM in LLMs. [tom3_kosinski2023theory, tom7_kosinski2024evaluating] report that GPT-3.5 and GPT-4 perform at levels matching or exceeding those of children in structured ToM tasks, suggesting that ToM-like behavior appears to develop as models scale in size. Through mechanistic interpretation, tom34_jamali2023unveiling offer evidence of parallels between the LLM embeddings and neurons in the human brain.
However, these claims have been challenged. tom1_ullman2023large, tom4_shapira2023clever, tom13_holterman2023does show that minor task changes often cause LLMs’ performance to drop, suggesting reliance on brittle heuristics rather than true semantic understanding. To address this, researchers have developed more robust benchmarks and evaluation protocols featuring procedurally generated, higher-order, or broader ToM tasks [tom5_gandhi2023understanding, tom14_chen2023tobench, tom12_he2023hi, tom32_xu2024opentom, tom41_shinoda2025tomato]. These benchmarks frequently reveal significant declines in performance on complex tasks, such as 6th-order belief attribution [tom9_street2024llms] or faux pas detection [tom6_strachan2024testing].
Some argue these failures stem from a lack of general commonsense reasoning rather than an inability to represent mental states [tom45_pi2024dissecting]. tom_riemer2024position_broken further suggest distinguishing between literal and functional ToM, noting that current benchmarks inadequately assess functional ToM. While newer models continue to improve on structured ToM tasks [tom6_strachan2024testing, tom9_street2024llms], they remain inconsistent in open-ended, adversarial, or pragmatic reasoning scenarios [tom22_sclar2024explore, tom27_yu2025persuasive, tom31_zhang2025tom]. Comparisons with human baselines show that LLMs can approximate ToM behavior in narrow contexts but still fall short of general human-like social cognition [tom6_strachan2024testing, tom11_van2023theory, tom20_jones2024comparing]. Recent work has also extended ToM evaluation to multimodal settings [tom37_chen2024mm, tom38_james2024gpt4o, tom42_jin2024mmtom] and multi-agent interactions [tom35_zhang2024tom].
@@
LLMs exhibit psychometrically measurable ToM-like reasoning, especially when appropriately prompted or structured, but current evidence suggests these capabilities depend on surface-level linguistic cues and lack robustness. The proficiency of LLMs in ToM remains a contentious issue. Interested readers are referred to recent comprehensive reviews on ToM in LLMs [saritacs2025tom_survey, tom_ma2023position, mao2024_machine_tom_review].
@@
EI is the subset of social intelligence that involves the ability to monitor one’s own and others’ feelings and emotions, to discriminate among them and to use this information to guide one’s thinking and actions [salovey1990emotional].
Recent work introduces benchmarks to assess EI in LLMs using structured, theory-driven tasks. Tools like EmoBench [sabour2024emobench] and EQ-Bench [paech2023eqbench] operationalize constructs such as emotional understanding and regulation, often referencing frameworks like Mayer-Salovey-Caruso Emotional Intelligence Test. Other studies adapt established psychometric instruments, including Situational Evaluation of Complex Emotional Understanding (SECEU) [wang2023emotional], the Levels of Emotional Awareness Scale (LEAS) [elyoseph2023chatgpt], the Toronto Alexithymia Scale (TAS-20) and Empathy Quotient (EQ-60) [patel2023identification], among others [huang2024apathetic]. Advanced models like GPT-4 often match or surpass human baselines in emotional awareness and understanding [wang2023emotional, elyoseph2023chatgpt, patel2023identification], though they lack deep reflexive analysis of emotional experience and motivation [vzorinab2024emotional]. Recent benchmarks also extend EI evaluation to multimodal settings [hu2025emobench].
@@
Advanced LLMs can perform on par with or better than humans in some tasks of EI [elyoseph2023chatgpt, patel2023identification]. However, they show evident limitations in several areas, such as displaying artificial or mechanical patterns when expressing empathy [lee2024large], the deep reflexive analysis of emotional experiences [vzorinab2024emotional], and the misalignment with human emotional behaviors [huang2023emotionbench].
@@
SI is the ability to understand and manage people [thorndike1937evaluation]. SI in LLMs determines how well these models interpret and respond to social situations. Psychometric tools applied to LLMs yield mixed results. si1_mittelstadt2024large use Situational Judgment Tests (SJTs) and find some LLMs outperform humans in expert-rated social appropriateness. However, LLMs still struggle with implicit social understanding: si2_shapira2023how show LLMs perform poorly on faux pas tests, and si11_xu2024academically find superficial friendliness often leads to errors in the Situational Evaluation of SI (SESI). The AgentSense benchmark further highlights LLMs’ limitations in complex social interactions, especially regarding high-level needs and private information [si8_mou2024agentsense].
Recent work expands to multi-agent environments. SOTOPIA [si3_zhou2023sotopia] and SOTOPIA-$\pi$ [si4_wang2024sotopia] simulate cooperative and competitive social contexts, while the STSS benchmark [si6_wang2024towards] evaluates SI through task-oriented simulations focused on outcomes and goals. The CogMir framework [si5_liu2024exploring] finds LLMs and humans are similarly consistent in irrational and prosocial decisions under uncertainty. DeSIQ [si12_guo2023desiq] extends SI evaluation to multimodal settings. Critically, si14_kovac2024socialai argue current benchmarks lack developmental grounding and propose the SocialAI school as a more comprehensive framework for studying LLM-based SI.
@@
LLMs show measurable competence in rule-based, socially appropriate behavior, particularly in structured environments. They perform well on: 1) following predefined social norms [si1_mittelstadt2024large]; 2) completing interactional goals in multi-agent settings [si3_zhou2023sotopia, si4_wang2024sotopia]; 3) replicating human-like prosocial decisions [si5_liu2024exploring]; and 4) being persuasive, even better than human persuaders [schoenegger2025large]. Weaker performance is noted in tasks where superficial friendliness causes errors [si11_xu2024academically], understanding high-level growth needs is essential [si8_mou2024agentsense], and LLMs must implicitly describe social situations [si2_shapira2023how].
@@ 4.2.3 Psychology of language
Psycholinguistics, a subfield of psychology, explores how humans comprehend, generate, and acquire language [carroll1986psychology]. Insights from this discipline help evaluate how LLMs process language and mirror human linguistic features.
@@
LLM language comprehension is evaluated across multiple linguistic levels—sound, word, syntax, meaning, and discourse. pl1_duan2024hlb introduce a benchmark of 10 psycholinguistic tests to assess these comprehensively. Many studies focus on specific aspects, such as syntactic and semantic processing [pl3_wang2024how, pl6_arehalli2022syntactic, pl8_wolfman2024hierarchical, pl5_wilcox2021targeted]. Surprisal, a measure of word predictability, is widely used to model processing difficulty in syntactic ambiguities and hierarchical structures [pl5_wilcox2021targeted, pl6_arehalli2022syntactic, pl17_li2024incremental]. Results are mixed: for example, GPT-2-XL shows human-like competence in sound-gender association and implicit causality, but not in sound-shape association [pl13_duan2024unveiling]. Other evaluation tasks include grammaticality judgment [pl18_qiu2024evaluating, pl29_ide2024how], pragmatic inference [pl28_bojic2023gpt, pl_hu2023fine, pl_ruis2023goldilocks], argument role processing [pl14_lee2024a], and discourse comprehension [pl1_duan2024hlb]. Advanced models like GPT-4 often match or surpass humans in pragmatic reasoning [pl28_bojic2023gpt] and grammaticality judgment [pl21_dentella2024language], though performance varies with prompt format [pl7_hu2023prompt] and implicature understanding remains limited [pl_ruis2023goldilocks].
@@
A rich line of research evaluates the creativity of LLM language generation [pl_stevenson2022putting, pl_hubert2023artificial, pl_chakrabarty2024art, pl_orwig2024language, pl_boussioux2024crowdless, bellemare2024divergent], using tests like Guilford’s Alternative Uses Test (AUT) [guilford1978alternate] and the Torrance Test of Creative Thinking (TTCT) [torrance1966torrance]. Early models such as GPT-3 lack originality and novelty, while advanced LLMs like GPT-4 surpass the human average in creativity. Notably, tang2024humanlike find LLMs underperform in divergent creativity (e.g., novel uses for familiar objects) but can rival humans in open-ended creative writing. cai2024large show that ChatGPT produces human-like responses in most of 12 psycholinguistic tests, but AI-generated analogies often lack human-like linguistic properties [pl_seals2023long]. LLM-generated stories tend to be positive and lack suspense or narrative diversity compared to human writing [pl_tian2024large]. Linguistic profiling has also been used to characterize the task-specific language abilities of LLMs [miaschi2024evaluating].
@@
Work on developmental plausibility examines whether LLMs replicate stages of language learning analogous to those in children [pl10_steuer2023large, shah2024development]. It is shown that regardless of model size, the developmental trajectories of pretrained language models consistently exhibit a window of maximal alignment with human cognitive development [shah2024development]. Another study by pl_frank2023bridging draws inspiration from human language development to explain the data inefficiency of LLMs; the relative efficiency of human language acquisition is possibly due to pre-existing conceptual knowledge, multimodal grounding, and the interactive, social nature of their input.
@@
Early foundations that evaluate language models based on BERT and LSTM link computational linguistics and psycholinguistic mechanisms [ettinger2020bert, futrell2019neural, arehalli2020neural]; these models generally fall short in diverse psycholinguistic tasks. Recent studies extend evaluations to LLMs and broader tasks [pl1_duan2024hlb], where advanced LLMs are shown to surpass humans on tasks such as pragmatic reasoning [pl28_bojic2023gpt] and creative writing [tang2024humanlike]. Some deficiencies still exist, such as limited implicature understanding [pl_ruis2023goldilocks], prompt-sensitive linguistic competence [pl7_hu2023prompt], and lack of human-like psycholinguistic properties [pl_seals2023long, pl_tian2024large]. In addition, alignment between LLMs and humans shows mixed results in both linguistic cognition [pl13_duan2024unveiling, pl8_wolfman2024hierarchical] as well as language acquisition [shah2024development, pl_frank2023bridging].